OpenStudy (anonymous):

complete the table' slope-intercept form standard form y =x+5 2x-3y=-6 y=6x-8 7x-y=9 y=-2x-4 2x+3y=-9 I need the slope-intercept form for the equation on the right and the standard form the equations on the left. SHOW WORK, plzzzz

OpenStudy (anonymous):

-x+y=5 y=2/3x-2 -6x+y=-8 y=7x+9 2x+y=-4 y=-3/2x-9